# Series Forecasting

### v2

@@ 5.2. Main Results
1) Compared with forecasters of different structures, DUET achieves an excellent predictive performance. From the perspective of absolute performance, DUET demonstrates a significant improvement against the second-best baseline PDF, with an impressive 6.9% reduction in MSE and a 6.2% reduction in MAE.
@@
3) DUET also demonstrates strong performance in tackling temporal heterogeneity caused by Temporal Distribution Shift. Compared to the state-of-the-art Non-stationary Transformer model for non-stationary time series modeling. DUET achieves a significant reduction of 32.3% in MSE and 21.5% in MAE. This result indicates that the proposed Temporal Clustering Module is more effective in modeling heterogeneous temporal patterns.

### v3

@@ 5.2. Main Results
1) Compared with forecasters of different structures, DUET achieves an excellent predictive performance. From the perspective of absolute performance, DUET demonstrates a significant improvement against the second-best baseline PDF, with an impressive 7.1% reduction in MSE and a 6.5% reduction in MAE.
@@
3) DUET also demonstrates strong performance in tackling temporal heterogeneity caused by Temporal Distribution Shift. Compared to the state-of-the-art Non-stationary Transformer model for non-stationary time series modeling. DUET achieves a significant reduction of 32.4% in MSE and 21.7% in MAE. This result indicates that the proposed Temporal Clustering Module is more effective in modeling heterogeneous temporal patterns.
@@ 6. Conclusions
Acknowledgements.
This work was partially supported by National Natural Science Foundation of China (62472174, 62372179). Jilin Hu is the corresponding author of the work.